Please help me to understand how I solve these types of questions, I'am having a lot of trouble with this
Solve for y
3y=ab
bx/y=r
ab=cy

Answer by ice_water35(8)

>>>>>
Solve for y
3y=ab
(3y)/3=(ab)/3
y=(ab)/3
bx/y=r
(bx/y)(y)=(r)(y)
bx=ry
bx/r=ry/r
bx/r=y
y=bx/r
ab=cy
ab/c=cy/c
ab/c=y
y=ab/c
If you have some further questions, you may contact me. this is my number: 09223011043